OpenStudy (anonymous):

Computer Craft pays its sales associates a 15% commission on every sale. Barry sold a laptop computer for $1,900 and a printer for $300. How much money did Barry earn on the sale of the items?

OpenStudy (anonymous):

15/100 * 1900 = 15/100 * 300 =

OpenStudy (anonymous):